# Cooperative Decentralized Backdoor Attacks on Vertical Federated Learning

### v1

@@ VI-C Varying Adversaries and Attack Network
Impact of Trigger Intensity: We next analyze the impact of the intensity value $\gamma$ on the ASR. The results across all three datasets are shown in 5(c). The analysis reveals that a stronger trigger produces a higher ASR, with the top-performing server model showing a stronger association to the target as the trigger’s intensity increases. Notably, while MNIST and CIFAR-10 achieve relatively good performance even at lower $\gamma$ values, Fashion-MNIST requires a higher trigger intensity to attain desirable ASR accuracy. This is likely due to a significant portion of Fashion-MNIST datapoints (being covered in white), matching the background of the trigger, thus necessitating a stronger trigger to differentiate from the clean features.
Graph Connectivity: Next, we investigate how the algebraic connectivity $\rho$ of the adversary graph affects the attack performance. We simulate this by progressively increasing the average degree of the graph, beginning from a ring topology, and consider Method 2 from Fig. 3 for trigger embedding. The results are given in Fig. 5(d). We see that the attack efficacy tends to get enhanced (i.e., ASR increases) with the increase of $\rho$, indicating that $\delta(\rho)$ from Theorem 1 is increasing in $\rho$. This is because the adversaries receive a higher share of the features for higher $\rho$. Note that due to the increased complexity of features in CIFAR10, the server becomes more reliant on the trigger, meaning a lower $\rho$ is sufficient to achieve a high ASR.
@@
Function Analysis of $\delta(\rho)$: Finally, we examine the relationship between $\delta(\rho)$ and $\rho$ from our analysis in Sec. V. We compute $\delta$ as the L2 norm difference between the gradients used to update the top model under attack and under benign conditions. To enable broader analysis of connectivity over a larger range of $\rho$, we resize the samples [55, 56, 57, 58] to $64\times 64$. This adjustment allows for accommodating 20 total clients, including 10 adversaries, thereby supporting a denser adversary graph. Starting with a ring topology, we incrementally add edges until the network becomes fully connected. Each adversary adopts a subtrigger size of $2\times 2$ with a total trigger area of $40$. For CIFAR-10 and SVHN, the VAE architecture is adjusted to a 2-layer encoder and decoder CVAE.
@@
Figure 6: The gradient perturbation presented in the top model from the adversaries for differing levels of connectivity $\rho$. This validates the hypothesis in Sec. V that $\delta(\rho)$ is an increasing function of $\rho$. As $\rho$ increases, the perturbation saturates quickly.
@@ VII Conclusion
Utilizing Assumption 1 and the above iterative equation, we have
@@
Summing the above inequality from $t=0$ to $T-1$ and utilizing the fact that the loss function is non-negative, i.e., $\mathbb{E}[F({\theta}^{T})]\geq 0$, we have

### v2

@@ VI-C Varying Adversaries and Attack Network
Impact of Trigger Intensity: We next analyze the impact of the intensity value $\gamma$ on the ASR. The results across three datasets are shown in 5(c). The analysis reveals that a stronger trigger produces a higher ASR, with the top-performing server model showing a stronger association to the target as the trigger’s intensity increases. Notably, while MNIST and CIFAR-10 achieve relatively good performance even at lower $\gamma$ values, Fashion-MNIST requires a higher trigger intensity to attain desirable ASR accuracy. This is likely due to a significant portion of Fashion-MNIST datapoints (being covered in white), matching the background of the trigger, thus necessitating a stronger trigger to differentiate from the clean features.
Graph Connectivity: Next, we investigate how the algebraic connectivity $\rho$ of the adversary graph affects the attack performance. We simulate this by progressively increasing the average degree of the graph, beginning from a line topology, and consider Method 2 from Fig. 3 for trigger embedding. The results are given in Fig. 5(d). We see that the attack efficacy tends to get enhanced (i.e., ASR increases) with the increase of $\rho$, indicating that $\delta(\rho)$ from Theorem 1 is increasing in $\rho$. This is because the adversaries receive a higher share of the features for higher $\rho$. Note that due to the increased complexity of features in CIFAR10, the server becomes more reliant on the trigger, meaning a lower $\rho$ is sufficient to achieve a high ASR.
@@
Function Analysis of $\delta(\rho)$: Finally, we examine the relationship between $\delta(\rho)$ and $\rho$ from our analysis in Sec. V. We compute $\delta$ as the L2 norm difference between the gradients used to update the top model under attack and under benign conditions. To enable broader analysis of connectivity over a larger range of $\rho$, we resize the samples [65, 66, 67, 68] to $64\times 64$. This adjustment allows for accommodating 20 total clients, including 10 adversaries, thereby supporting a denser adversary graph. Starting with a line topology, we incrementally add edges until the network becomes fully connected. Each adversary adopts a subtrigger size of $2\times 2$ with a total trigger area of $40$. For CIFAR-10 and SVHN, the VAE architecture is adjusted to a 2-layer encoder and decoder CVAE.
@@
Figure 6: The gradient perturbation in the top model from the adversaries for differing levels of connectivity $\rho$. This validates the hypothesis in Sec. V that $\delta(\rho)$ is a non-decreasing function of $\rho$. As $\rho$ increases, the perturbation saturates quickly. When no perturbation takes place, it is marked with an “x”, as seen for some datasets when $\rho\in\{0.1,0.6$}.
@@ VII Conclusion
From Assumption 1 and the above iterative equation, we have
@@
Summing the above inequality from $t=0$ to T1T-1 and utilizing the fact that the loss function is non-negative, i.e., $\mathbb{E}[F({\theta}^{T})]\geq 0$, we have